OpenStudy (anonymous):

double checking: v=y^-3 :: y=v^-1/3 and what's a good way to remember how to solve these things?

OpenStudy (jdoe0001):

hmm... what's there to solve?

OpenStudy (anonymous):

\[v=y ^{-3}\] \[y=v ^{-1/3}\] and do the threes in those equations look like twos to anyone else?

OpenStudy (anonymous):

no real solving...is that right?

OpenStudy (jdoe0001):

they look good to me, nothing to solve, the equalities are proper

OpenStudy (anonymous):

I guess I have an aversion to fractional exponents - I just can never convince myself they are right

OpenStudy (jdoe0001):

hehhe well, they're very useful for manipulations just keep in mind that \(\bf \large {a^{\frac{n}{m}} = \sqrt[m]{a^n}\\\quad \\ a^{-\frac{n}{m}} = \cfrac{1}{a^{\frac{n}{m}}} \implies \cfrac{1}{\sqrt[m]{a^n}}\\\quad \\ a^{-n} = \cfrac{1}{a^n}}\)